# Josephson diodes induced by the loop current states

### v1

@@ Abstract
Discussion.— In summary, we have demonstrated that the JDE can be realized in the JJs constructed by the monolayer and bilayer honeycomb systems with loop current order, which directly breaks the time-reversal symmetry. For the monolayer system, the inversion symmetry can be broken by either applying a staggered sublattice potential for the HM or by reverting parts of the loop current directions to form the MHM, leading to the JDE. For the bilayer system, the JDE can be realized by applying an external electric field perpendicular to the system resulting in an electrically controllable Josephson diode, or by stacking the two layers of HMs with opposite loop current directions, which leads to an intrinsic field-free Josephson diode. Moreover, we also find that the JDE disappears when the JJ is constructed along the armchair edge which has an extra symmetry to prevent the emergence of the JDE.
@@
We expect this mechanism to realize the JDE can be easily applied to the other hexagonal system such as the Kagome systems. One of the possible platforms to realize this field-free JDE induced by the loop current is the Kagome metal system, which hosts a special charge density wave state as a promising candidate for the loop current order [84, 85, 88]. We hope it stimulates the exploration of the field-free JDE and expands the scope of the possible material systems to realize this effect.
@@ Acknowledgements.
This work is supported by National Natural Science Foundation of China (NSFC) Grants No. 12274279. We gratefully acknowledge HZWTECH for providing computation facilities.References
- Jiang and Hu [2022] K. Jiang and J. Hu, Superconducting diode effects, Nature Physics 18, 1145 (2022).
@@
- Nadeem et al. [2023] M. Nadeem, M. S. Fuhrer, and X. Wang, The superconducting diode effect, Nature Reviews Physics 5, 558 (2023).
@@
- Ando et al. [2020] F. Ando, Y. Miyasaka, T. Li, J. Ishizuka, T. Arakawa, Y. Shiota, T. Moriyama, Y. Yanase, and T. Ono, Observation of superconducting diode effect, Nature 584, 373 (2020).
@@
- Bauriedl et al. [2022] L. Bauriedl, C. Bäuml, L. Fuchs, C. Baumgartner, N. Paulik, J. M. Bauer, K.-Q. Lin, J. M. Lupton, T. Taniguchi, K. Watanabe, C. Strunk, and N. Paradiso, Supercurrent diode effect and magnetochiral anisotropy in few-layer nbse2, Nature Communications 13, 4266 (2022).
@@
- Yun et al. [2023] J. Yun, S. Son, J. Shin, G. Park, K. Zhang, Y. J. Shin, J.-G. Park, and D. Kim, Magnetic proximity-induced superconducting diode effect and infinite magnetoresistance in a van der waals heterostructure, Phys. Rev. Res. 5, L022064 (2023).
@@
- Lin et al. [2022] J.-X. Lin, P. Siriviboon, H. D. Scammell, S. Liu, D. Rhodes, K. Watanabe, T. Taniguchi, J. Hone, M. S. Scheurer, and J. I. A. Li, Zero-field superconducting diode effect in small-twist-angle trilayer graphene, Nature Physics 18, 1221 (2022).
@@
- Le et al. [2024] T. Le, Z. Pan, Z. Xu, J. Liu, J. Wang, Z. Lou, X. Yang, Z. Wang, Y. Yao, C. Wu, and X. Lin, Superconducting diode effect and interference patterns in kagome csv3sb5, Nature 630, 64 (2024).
@@
- Wakatsuki et al. [2017] R. Wakatsuki, Y. Saito, S. Hoshino, Y. M. Itahashi, T. Ideue, M. Ezawa, Y. Iwasa, and N. Nagaosa, Nonreciprocal charge transport in noncentrosymmetric superconductors, Science Advances 3, e1602390 (2017).
@@
- Miyasaka et al. [2021] Y. Miyasaka, R. Kawarazaki, H. Narita, F. Ando, Y. Ikeda, R. Hisatomi, A. Daido, Y. Shiota, T. Moriyama, Y. Yanase, and T. Ono, Observation of nonreciprocal superconducting critical field, Applied Physics Express 14, 073003 (2021).
@@
- Zhang et al. [2020] E. Zhang, X. Xu, Y.-C. Zou, L. Ai, X. Dong, C. Huang, P. Leng, S. Liu, Y. Zhang, Z. Jia, X. Peng, M. Zhao, Y. Yang, Z. Li, H. Guo, S. J. Haigh, N. Nagaosa, J. Shen, and F. Xiu, Nonreciprocal superconducting nbse2 antenna, Nature Communications 11, 5634 (2020).
@@
- Lyu et al. [2021] Y.-Y. Lyu, J. Jiang, Y.-L. Wang, Z.-L. Xiao, S. Dong, Q.-H. Chen, M. V. Milošević, H. Wang, R. Divan, J. E. Pearson, P. Wu, F. M. Peeters, and W.-K. Kwok, Superconducting diode effect via conformal-mapped nanoholes, Nature Communications 12, 2703 (2021).
@@
- Narita et al. [2022] H. Narita, J. Ishizuka, R. Kawarazaki, D. Kan, Y. Shiota, T. Moriyama, Y. Shimakawa, A. V. Ognev, A. S. Samardak, Y. Yanase, and T. Ono, Field-free superconducting diode effect in noncentrosymmetric superconductor/ferromagnet multilayers, Nature Nanotechnology 17, 823 (2022).
@@
- Strambini et al. [2022] E. Strambini, M. Spies, N. Ligato, S. Ilić, M. Rouco, C. González-Orellana, M. Ilyn, C. Rogero, F. S. Bergeret, J. S. Moodera, P. Virtanen, T. T. Heikkilä, and F. Giazotto, Superconducting spintronic tunnel diode, Nature Communications 13, 2431 (2022).
@@
- Wu et al. [2022] H. Wu, Y. Wang, Y. Xu, P. K. Sivakumar, C. Pasco, U. Filippozzi, S. S. P. Parkin, Y.-J. Zeng, T. McQueen, and M. N. Ali, The field-free josephson diode in a van der waals heterostructure, Nature 604, 653 (2022).
@@
- Baumgartner et al. [2022a] C. Baumgartner, L. Fuchs, A. Costa, S. Reinhardt, S. Gronin, G. C. Gardner, T. Lindemann, M. J. Manfra, P. E. Faria Junior, D. Kochan, J. Fabian, N. Paradiso, and C. Strunk, Supercurrent rectification and magnetochiral effects in symmetric josephson junctions, Nature Nanotechnology 17, 39 (2022a).
@@
- Baumgartner et al. [2022b] C. Baumgartner, L. Fuchs, A. Costa, J. Picó-Cortés, S. Reinhardt, S. Gronin, G. C. Gardner, T. Lindemann, M. J. Manfra, P. E. F. Junior, D. Kochan, J. Fabian, N. Paradiso, and C. Strunk, Effect of rashba and dresselhaus spin–orbit coupling on supercurrent rectification and magnetochiral anisotropy of ballistic josephson junctions, Journal of Physics: Condensed Matter 34, 154005 (2022b).
@@
- Pal et al. [2022] B. Pal, A. Chakraborty, P. K. Sivakumar, M. Davydova, A. K. Gopi, A. K. Pandeya, J. A. Krieger, Y. Zhang, M. Date, S. Ju, N. Yuan, N. B. M. Schröter, L. Fu, and S. S. P. Parkin, Josephson diode effect from cooper pair momentum in a topological semimetal, Nature Physics 18, 1228 (2022).
@@
- Díez-Mérida et al. [2023] J. Díez-Mérida, A. Díez-Carlón, S. Y. Yang, Y. M. Xie, X. J. Gao, J. Senior, K. Watanabe, T. Taniguchi, X. Lu, A. P. Higginbotham, K. T. Law, and D. K. Efetov, Symmetry-broken josephson junctions and superconducting diodes in magic-angle twisted bilayer graphene, Nature Communications 14, 2396 (2023).
@@
- Gupta et al. [2023] M. Gupta, G. V. Graziano, M. Pendharkar, J. T. Dong, C. P. Dempsey, C. Palmstrøm, and V. S. Pribiag, Gate-tunable superconducting diode effect in a three-terminal josephson device, Nature Communications 14, 3078 (2023).
@@
- Turini et al. [2022] B. Turini, S. Salimian, M. Carrega, A. Iorio, E. Strambini, F. Giazotto, V. Zannier, L. Sorba, and S. Heun, Josephson diode effect in high-mobility insb nanoflags, Nano Letters 22, 8502 (2022).
@@
- Chen et al. [2024a] P. Chen, G. Wang, B. Ye, J. Wang, L. Zhou, Z. Tang, L. Wang, J. Wang, W. Zhang, J. Mei, W. Chen, and H. He, Edelstein Effect Induced Superconducting Diode Effect in Inversion Symmetry Breaking MoTe2 Josephson Junctions, Advanced Functional Materials 34, 2311229 (2024a).
@@
- Ghosh et al. [2024] S. Ghosh, V. Patil, A. Basu, Kuldeep, A. Dutta, D. A. Jangade, R. Kulkarni, A. Thamizhavel, J. F. Steiner, F. von Oppen, and M. M. Deshmukh, High-temperature Josephson diode, Nature Materials 23, 612 (2024).
@@
- Lotfizadeh et al. [2024] N. Lotfizadeh, W. F. Schiela, B. Pekerten, P. Yu, B. H. Elfeky, W. M. Strickland, A. Matos-Abiague, and J. Shabani, Superconducting diode effect sign change in epitaxial Al-InAs Josephson junctions, Communications Physics 7, 1 (2024).
@@
- Kim et al. [2024] J.-K. Kim, K.-R. Jeon, P. K. Sivakumar, J. Jeon, C. Koerner, G. Woltersdorf, and S. S. P. Parkin, Intrinsic supercurrent non-reciprocity coupled to the crystal structure of a van der Waals Josephson barrier, Nature Communications 15, 1120 (2024).
- Hu et al. [2007] J. Hu, C. Wu, and X. Dai, Proposed design of a josephson diode, Phys. Rev. Lett. 99, 067004 (2007).
@@
- Buzdin [2008] A. Buzdin, Direct coupling between magnetism and superconducting current in the josephson junction, Phys. Rev. Lett. 101, 107005 (2008).
@@
- Huang and Hu [2014] Z. Huang and X. Hu, Josephson effects in three-band superconductors with broken time-reversal symmetry, Applied Physics Letters 104, 162602 (2014).
@@
- Yokoyama et al. [2014] T. Yokoyama, M. Eto, and Y. V. Nazarov, Anomalous josephson effect induced by spin-orbit interaction and zeeman effect in semiconductor nanowires, Phys. Rev. B 89, 195407 (2014).
- Dolcini et al. [2015] F. Dolcini, M. Houzet, and J. S. Meyer, Topological josephson junctions, Phys. Rev. B 92, 035428 (2015).
@@
- Szombati et al. [2016] D. B. Szombati, S. Nadj-Perge, D. Car, S. R. Plissard, E. P. A. M. Bakkers, and L. P. Kouwenhoven, Josephson 0-junction in nanowire quantum dots, Nature Physics 12, 568 (2016).
@@
- Assouline et al. [2019] A. Assouline, C. Feuillet-Palma, N. Bergeal, T. Zhang, A. Mottaghizadeh, A. Zimmers, E. Lhuillier, M. Eddrie, P. Atkinson, M. Aprili, and H. Aubin, Spin-orbit induced phase-shift in bi2se3 josephson junctions, Nature Communications 10, 126 (2019).
@@
- Alidoust and Linder [2013] M. Alidoust and J. Linder, -state and inverted fraunhofer pattern in nonaligned josephson junctions, Phys. Rev. B 87, 060503 (2013).
@@
- Bobkova et al. [2016] I. V. Bobkova, A. M. Bobkov, A. A. Zyuzin, and M. Alidoust, Magnetoelectrics in disordered topological insulator josephson junctions, Phys. Rev. B 94, 134506 (2016).
@@
- Alidoust and Hamzehpour [2017] M. Alidoust and H. Hamzehpour, Spontaneous supercurrent and phase shift parallel to magnetized topological insulator interfaces, Phys. Rev. B 96, 165422 (2017).
@@
- Alidoust et al. [2018] M. Alidoust, M. Willatzen, and A.-P. Jauho, Strain-engineered majorana zero energy modes and josephson state in black phosphorus, Phys. Rev. B 98, 085414 (2018).
@@
- Alidoust [2020] M. Alidoust, Critical supercurrent and state for probing a persistent spin helix, Phys. Rev. B 101, 155123 (2020).
@@
- Alidoust et al. [2021] M. Alidoust, C. Shen, and I. Žutić, Cubic spin-orbit coupling and anomalous josephson effect in planar junctions, Phys. Rev. B 103, L060503 (2021).
@@
- Halterman et al. [2022] K. Halterman, M. Alidoust, R. Smith, and S. Starr, Supercurrent diode effect, spin torques, and robust zero-energy peak in planar half-metallic trilayers, Phys. Rev. B 105, 104508 (2022).
- Misaki and Nagaosa [2021] K. Misaki and N. Nagaosa, Theory of the nonreciprocal josephson effect, Phys. Rev. B 103, 245302 (2021).
@@
- He et al. [2022] J. J. He, Y. Tanaka, and N. Nagaosa, A phenomenological theory of superconductor diodes, New Journal of Physics 24, 053014 (2022).
@@
- Yuan and Fu [2022] N. F. Q. Yuan and L. Fu, Supercurrent diode effect and finite-momentum superconductors, Proceedings of the National Academy of Sciences 119, e2119548119 (2022).
@@
- Daido et al. [2022] A. Daido, Y. Ikeda, and Y. Yanase, Intrinsic superconducting diode effect, Phys. Rev. Lett. 128, 037001 (2022).
@@
- Davydova et al. [2022] M. Davydova, S. Prembabu, and L. Fu, Universal josephson diode effect, Science Advances 8, eabo0309 (2022).
@@
- Scammell et al. [2022] H. D. Scammell, J. I. A. Li, and M. S. Scheurer, Theory of zero-field superconducting diode effect in twisted trilayer graphene, 2D Materials 9, 025027 (2022).
@@
- Zhang et al. [2022] Y. Zhang, Y. Gu, P. Li, J. Hu, and K. Jiang, General Theory of Josephson Diodes, Physical Review X 12, 041013 (2022).
@@
- Souto et al. [2022] R. S. Souto, M. Leijnse, and C. Schrade, Josephson diode effect in supercurrent interferometers, Phys. Rev. Lett. 129, 267702 (2022).
@@
- Karabassov et al. [2022] T. Karabassov, I. V. Bobkova, A. A. Golubov, and A. S. Vasenko, Hybrid helical state and superconducting diode effect in superconductor/ferromagnet/topological insulator heterostructures, Phys. Rev. B 106, 224509 (2022).

### v3

@@ VIII Discussion
In summary, we have demonstrated that the JDE can be realized in the JJs constructed by the monolayer and bilayer honeycomb systems with loop current order, which directly breaks the time-reversal symmetry. For the monolayer system, the inversion symmetry can be broken by either applying a staggered sublattice potential for the HM or by reverting parts of the loop current directions to form the MHM, leading to the JDE. For the bilayer system, the JDE can be realized by applying an external electric field perpendicular to the system resulting in an electrically controllable Josephson diode, or by stacking the two layers of HMs with opposite loop current directions, which leads to an intrinsic field-free Josephson diode. Moreover, we also find that the JDE disappears when the JJ is constructed along the armchair edge which has an extra symmetry to prevent the emergence of the JDE. Finally, we demonstrate that the JDE is robust against disorder induced by impurity by calculating the average CPR for the systems with random onsite potential. The JDE persists in the average CPR over 100 random configurations. While this work focuses on the short-junction limit, we expect the JDE to persist in longer junctions, since the symmetry-breaking pattern remains unchanged as the junction size increases.
@@
We expect this mechanism to realize the JDE can be easily applied to other hexagonal systems such as the Kagome systems. One of the possible platforms to realize this field-free JDE induced by the loop current is the Kagome metal system, which hosts a special charge density wave state as a promising candidate for the loop current orderΒ [89, 90, 93]. We hope it stimulates the exploration of the field-free JDE and expands the scope of the possible material systems to realize this effect.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (NSFC) Grants No. 12274279. We gratefully acknowledge HZWTECH for providing computation facilities.References
- JiangΒ andΒ Hu [2022] K.Β JiangΒ andΒ J.Β Hu,Β Superconducting diode effects,Β Nature PhysicsΒ 18,Β 1145 (2022).
@@
- NadeemΒ etΒ al. [2023] M.Β Nadeem, M.Β S.Β Fuhrer,Β andΒ X.Β Wang,Β The superconducting diode effect,Β Nature Reviews PhysicsΒ 5,Β 558 (2023).
@@
- AndoΒ etΒ al. [2020] F.Β Ando, Y.Β Miyasaka, T.Β Li, J.Β Ishizuka, T.Β Arakawa, Y.Β Shiota, T.Β Moriyama, Y.Β Yanase,Β andΒ T.Β Ono,Β Observation of superconducting diode effect,Β NatureΒ 584,Β 373 (2020).
@@
- BauriedlΒ etΒ al. [2022] L.Β Bauriedl, C.Β BΓ€uml, L.Β Fuchs, C.Β Baumgartner, N.Β Paulik, J.Β M.Β Bauer, K.-Q.Β Lin, J.Β M.Β Lupton, T.Β Taniguchi, K.Β Watanabe, C.Β Strunk,Β andΒ N.Β Paradiso,Β Supercurrent diode effect and magnetochiral anisotropy in few-layer nbse2,Β Nature CommunicationsΒ 13,Β 4266 (2022).
@@
- YunΒ etΒ al. [2023] J.Β Yun, S.Β Son, J.Β Shin, G.Β Park, K.Β Zhang, Y.Β J.Β Shin, J.-G.Β Park,Β andΒ D.Β Kim,Β Magnetic proximity-induced superconducting diode effect and infinite magnetoresistance in a van der waals heterostructure,Β Phys. Rev. Res.Β 5,Β L022064 (2023).
@@
- LinΒ etΒ al. [2022] J.-X.Β Lin, P.Β Siriviboon, H.Β D.Β Scammell, S.Β Liu, D.Β Rhodes, K.Β Watanabe, T.Β Taniguchi, J.Β Hone, M.Β S.Β Scheurer,Β andΒ J.Β I.Β A.Β Li,Β Zero-field superconducting diode effect in small-twist-angle trilayer graphene,Β Nature PhysicsΒ 18,Β 1221 (2022).
@@
- LeΒ etΒ al. [2024] T.Β Le, Z.Β Pan, Z.Β Xu, J.Β Liu, J.Β Wang, Z.Β Lou, X.Β Yang, Z.Β Wang, Y.Β Yao, C.Β Wu,Β andΒ X.Β Lin,Β Superconducting diode effect and interference patterns in kagome csv3sb5,Β NatureΒ 630,Β 64 (2024).
@@
- WakatsukiΒ etΒ al. [2017] R.Β Wakatsuki, Y.Β Saito, S.Β Hoshino, Y.Β M.Β Itahashi, T.Β Ideue, M.Β Ezawa, Y.Β Iwasa,Β andΒ N.Β Nagaosa,Β Nonreciprocal charge transport in noncentrosymmetric superconductors,Β Science AdvancesΒ 3,Β e1602390 (2017).
@@
- MiyasakaΒ etΒ al. [2021] Y.Β Miyasaka, R.Β Kawarazaki, H.Β Narita, F.Β Ando, Y.Β Ikeda, R.Β Hisatomi, A.Β Daido, Y.Β Shiota, T.Β Moriyama, Y.Β Yanase,Β andΒ T.Β Ono,Β Observation of nonreciprocal superconducting critical field,Β Applied Physics ExpressΒ 14,Β 073003 (2021).
@@
- ZhangΒ etΒ al. [2020] E.Β Zhang, X.Β Xu, Y.-C.Β Zou, L.Β Ai, X.Β Dong, C.Β Huang, P.Β Leng, S.Β Liu, Y.Β Zhang, Z.Β Jia, X.Β Peng, M.Β Zhao, Y.Β Yang, Z.Β Li, H.Β Guo, S.Β J.Β Haigh, N.Β Nagaosa, J.Β Shen,Β andΒ F.Β Xiu,Β Nonreciprocal superconducting nbse2 antenna,Β Nature CommunicationsΒ 11,Β 5634 (2020).
@@
- LyuΒ etΒ al. [2021] Y.-Y.Β Lyu, J.Β Jiang, Y.-L.Β Wang, Z.-L.Β Xiao, S.Β Dong, Q.-H.Β Chen, M.Β V.Β MiloΕ‘eviΔ, H.Β Wang, R.Β Divan, J.Β E.Β Pearson, P.Β Wu, F.Β M.Β Peeters,Β andΒ W.-K.Β Kwok,Β Superconducting diode effect via conformal-mapped nanoholes,Β Nature CommunicationsΒ 12,Β 2703 (2021).
@@
- NaritaΒ etΒ al. [2022] H.Β Narita, J.Β Ishizuka, R.Β Kawarazaki, D.Β Kan, Y.Β Shiota, T.Β Moriyama, Y.Β Shimakawa, A.Β V.Β Ognev, A.Β S.Β Samardak, Y.Β Yanase,Β andΒ T.Β Ono,Β Field-free superconducting diode effect in noncentrosymmetric superconductor/ferromagnet multilayers,Β Nature NanotechnologyΒ 17,Β 823 (2022).
@@
- StrambiniΒ etΒ al. [2022] E.Β Strambini, M.Β Spies, N.Β Ligato, S.Β IliΔ, M.Β Rouco, C.Β GonzΓ‘lez-Orellana, M.Β Ilyn, C.Β Rogero, F.Β S.Β Bergeret, J.Β S.Β Moodera, P.Β Virtanen, T.Β T.Β HeikkilΓ€,Β andΒ F.Β Giazotto,Β Superconducting spintronic tunnel diode,Β Nature CommunicationsΒ 13,Β 2431 (2022).
@@
- WuΒ etΒ al. [2022] H.Β Wu, Y.Β Wang, Y.Β Xu, P.Β K.Β Sivakumar, C.Β Pasco, U.Β Filippozzi, S.Β S.Β P.Β Parkin, Y.-J.Β Zeng, T.Β McQueen,Β andΒ M.Β N.Β Ali,Β The field-free josephson diode in a van der waals heterostructure,Β NatureΒ 604,Β 653 (2022).
@@
- BaumgartnerΒ etΒ al. [2022a] C.Β Baumgartner, L.Β Fuchs, A.Β Costa, S.Β Reinhardt, S.Β Gronin, G.Β C.Β Gardner, T.Β Lindemann, M.Β J.Β Manfra, P.Β E.Β FariaΒ Junior, D.Β Kochan, J.Β Fabian, N.Β Paradiso,Β andΒ C.Β Strunk,Β Supercurrent rectification and magnetochiral effects in symmetric josephson junctions,Β Nature NanotechnologyΒ 17,Β 39 (2022a).
@@
- BaumgartnerΒ etΒ al. [2022b] C.Β Baumgartner, L.Β Fuchs, A.Β Costa, J.Β PicΓ³-CortΓ©s, S.Β Reinhardt, S.Β Gronin, G.Β C.Β Gardner, T.Β Lindemann, M.Β J.Β Manfra, P.Β E.Β F.Β Junior, D.Β Kochan, J.Β Fabian, N.Β Paradiso,Β andΒ C.Β Strunk,Β Effect of rashba and dresselhaus spinβorbit coupling on supercurrent rectification and magnetochiral anisotropy of ballistic josephson junctions,Β Journal of Physics: Condensed MatterΒ 34,Β 154005 (2022b).
@@
- PalΒ etΒ al. [2022] B.Β Pal, A.Β Chakraborty, P.Β K.Β Sivakumar, M.Β Davydova, A.Β K.Β Gopi, A.Β K.Β Pandeya, J.Β A.Β Krieger, Y.Β Zhang, M.Β Date, S.Β Ju, N.Β Yuan, N.Β B.Β M.Β SchrΓΆter, L.Β Fu,Β andΒ S.Β S.Β P.Β Parkin,Β Josephson diode effect from cooper pair momentum in a topological semimetal,Β Nature PhysicsΒ 18,Β 1228 (2022).
@@
- DΓez-MΓ©ridaΒ etΒ al. [2023] J.Β DΓez-MΓ©rida, A.Β DΓez-CarlΓ³n, S.Β Y.Β Yang, Y.Β M.Β Xie, X.Β J.Β Gao, J.Β Senior, K.Β Watanabe, T.Β Taniguchi, X.Β Lu, A.Β P.Β Higginbotham, K.Β T.Β Law,Β andΒ D.Β K.Β Efetov,Β Symmetry-broken josephson junctions and superconducting diodes in magic-angle twisted bilayer graphene,Β Nature CommunicationsΒ 14,Β 2396 (2023).
@@
- GuptaΒ etΒ al. [2023] M.Β Gupta, G.Β V.Β Graziano, M.Β Pendharkar, J.Β T.Β Dong, C.Β P.Β Dempsey, C.Β PalmstrΓΈm,Β andΒ V.Β S.Β Pribiag,Β Gate-tunable superconducting diode effect in a three-terminal josephson device,Β Nature CommunicationsΒ 14,Β 3078 (2023).
@@
- TuriniΒ etΒ al. [2022] B.Β Turini, S.Β Salimian, M.Β Carrega, A.Β Iorio, E.Β Strambini, F.Β Giazotto, V.Β Zannier, L.Β Sorba,Β andΒ S.Β Heun,Β Josephson diode effect in high-mobility insb nanoflags,Β Nano LettersΒ 22,Β 8502 (2022).
@@
- ChenΒ etΒ al. [2024a] P.Β Chen, G.Β Wang, B.Β Ye, J.Β Wang, L.Β Zhou, Z.Β Tang, L.Β Wang, J.Β Wang, W.Β Zhang, J.Β Mei, W.Β Chen,Β andΒ H.Β He,Β Edelstein Effect Induced Superconducting Diode Effect in Inversion Symmetry Breaking MoTe2 Josephson Junctions,Β Advanced Functional MaterialsΒ 34,Β 2311229 (2024a).
@@
- GhoshΒ etΒ al. [2024] S.Β Ghosh, V.Β Patil, A.Β Basu, Kuldeep, A.Β Dutta, D.Β A.Β Jangade, R.Β Kulkarni, A.Β Thamizhavel, J.Β F.Β Steiner, F.Β von Oppen,Β andΒ M.Β M.Β Deshmukh,Β High-temperature Josephson diode,Β Nature MaterialsΒ 23,Β 612 (2024).
@@
- LotfizadehΒ etΒ al. [2024] N.Β Lotfizadeh, W.Β F.Β Schiela, B.Β Pekerten, P.Β Yu, B.Β H.Β Elfeky, W.Β M.Β Strickland, A.Β Matos-Abiague,Β andΒ J.Β Shabani,Β Superconducting diode effect sign change in epitaxial Al-InAs Josephson junctions,Β Communications PhysicsΒ 7,Β 1 (2024).
@@
- KimΒ etΒ al. [2024] J.-K.Β Kim, K.-R.Β Jeon, P.Β K.Β Sivakumar, J.Β Jeon, C.Β Koerner, G.Β Woltersdorf,Β andΒ S.Β S.Β P.Β Parkin,Β Intrinsic supercurrent non-reciprocity coupled to the crystal structure of a van der Waals Josephson barrier,Β Nature CommunicationsΒ 15,Β 1120 (2024).
- HuΒ etΒ al. [2007] J.Β Hu, C.Β Wu,Β andΒ X.Β Dai,Β Proposed design of a josephson diode,Β Phys. Rev. Lett.Β 99,Β 067004 (2007).
@@
- Buzdin [2008] A.Β Buzdin,Β Direct coupling between magnetism and superconducting current in the josephson junction,Β Phys. Rev. Lett.Β 101,Β 107005 (2008).
@@
- HuangΒ andΒ Hu [2014] Z.Β HuangΒ andΒ X.Β Hu,Β Josephson effects in three-band superconductors with broken time-reversal symmetry,Β Applied Physics LettersΒ 104,Β 162602 (2014).
@@
- YokoyamaΒ etΒ al. [2014] T.Β Yokoyama, M.Β Eto,Β andΒ Y.Β V.Β Nazarov,Β Anomalous josephson effect induced by spin-orbit interaction and zeeman effect in semiconductor nanowires,Β Phys. Rev. BΒ 89,Β 195407 (2014).
- DolciniΒ etΒ al. [2015] F.Β Dolcini, M.Β Houzet,Β andΒ J.Β S.Β Meyer,Β Topological josephson junctions,Β Phys. Rev. BΒ 92,Β 035428 (2015).
@@
- SzombatiΒ etΒ al. [2016] D.Β B.Β Szombati, S.Β Nadj-Perge, D.Β Car, S.Β R.Β Plissard, E.Β P. A.Β M.Β Bakkers,Β andΒ L.Β P.Β Kouwenhoven,Β Josephson 0-junction in nanowire quantum dots,Β Nature PhysicsΒ 12,Β 568 (2016).
@@
- AssoulineΒ etΒ al. [2019] A.Β Assouline, C.Β Feuillet-Palma, N.Β Bergeal, T.Β Zhang, A.Β Mottaghizadeh, A.Β Zimmers, E.Β Lhuillier, M.Β Eddrie, P.Β Atkinson, M.Β Aprili,Β andΒ H.Β Aubin,Β Spin-orbit induced phase-shift in bi2se3 josephson junctions,Β Nature CommunicationsΒ 10,Β 126 (2019).
@@
- AlidoustΒ andΒ Linder [2013] M.Β AlidoustΒ andΒ J.Β Linder,Β -state and inverted fraunhofer pattern in nonaligned josephson junctions,Β Phys. Rev. BΒ 87,Β 060503 (2013).
@@
- BobkovaΒ etΒ al. [2016] I.Β V.Β Bobkova, A.Β M.Β Bobkov, A.Β A.Β Zyuzin,Β andΒ M.Β Alidoust,Β Magnetoelectrics in disordered topological insulator josephson junctions,Β Phys. Rev. BΒ 94,Β 134506 (2016).
@@
- AlidoustΒ andΒ Hamzehpour [2017] M.Β AlidoustΒ andΒ H.Β Hamzehpour,Β Spontaneous supercurrent and phase shift parallel to magnetized topological insulator interfaces,Β Phys. Rev. BΒ 96,Β 165422 (2017).
@@
- AlidoustΒ etΒ al. [2018] M.Β Alidoust, M.Β Willatzen,Β andΒ A.-P.Β Jauho,Β Strain-engineered majorana zero energy modes and josephson state in black phosphorus,Β Phys. Rev. BΒ 98,Β 085414 (2018).
@@
- Alidoust [2020] M.Β Alidoust,Β Critical supercurrent and state for probing a persistent spin helix,Β Phys. Rev. BΒ 101,Β 155123 (2020).
@@
- AlidoustΒ etΒ al. [2021] M.Β Alidoust, C.Β Shen,Β andΒ I.Β Ε½utiΔ,Β Cubic spin-orbit coupling and anomalous josephson effect in planar junctions,Β Phys. Rev. BΒ 103,Β L060503 (2021).
@@
- HaltermanΒ etΒ al. [2022] K.Β Halterman, M.Β Alidoust, R.Β Smith,Β andΒ S.Β Starr,Β Supercurrent diode effect, spin torques, and robust zero-energy peak in planar half-metallic trilayers,Β Phys. Rev. BΒ 105,Β 104508 (2022).
- MisakiΒ andΒ Nagaosa [2021] K.Β MisakiΒ andΒ N.Β Nagaosa,Β Theory of the nonreciprocal josephson effect,Β Phys. Rev. BΒ 103,Β 245302 (2021).
@@
- HeΒ etΒ al. [2022] J.Β J.Β He, Y.Β Tanaka,Β andΒ N.Β Nagaosa,Β A phenomenological theory of superconductor diodes,Β New Journal of PhysicsΒ 24,Β 053014 (2022).
@@
- YuanΒ andΒ Fu [2022] N.Β F.Β Q.Β YuanΒ andΒ L.Β Fu,Β Supercurrent diode effect and finite-momentum superconductors,Β Proceedings of the National Academy of SciencesΒ 119,Β e2119548119 (2022).
@@
- DaidoΒ etΒ al. [2022] A.Β Daido, Y.Β Ikeda,Β andΒ Y.Β Yanase,Β Intrinsic superconducting diode effect,Β Phys. Rev. Lett.Β 128,Β 037001 (2022).
@@
- DavydovaΒ etΒ al. [2022] M.Β Davydova, S.Β Prembabu,Β andΒ L.Β Fu,Β Universal josephson diode effect,Β Science AdvancesΒ 8,Β eabo0309 (2022).
@@
- ScammellΒ etΒ al. [2022] H.Β D.Β Scammell, J.Β I.Β A.Β Li,Β andΒ M.Β S.Β Scheurer,Β Theory of zero-field superconducting diode effect in twisted trilayer graphene,Β 2D MaterialsΒ 9,Β 025027 (2022).
@@
- ZhangΒ etΒ al. [2022] Y.Β Zhang, Y.Β Gu, P.Β Li, J.Β Hu,Β andΒ K.Β Jiang,Β General Theory of Josephson Diodes,Β Physical Review XΒ 12,Β 041013 (2022).
@@
- SoutoΒ etΒ al. [2022] R.Β S.Β Souto, M.Β Leijnse,Β andΒ C.Β Schrade,Β Josephson diode effect in supercurrent interferometers,Β Phys. Rev. Lett.Β 129,Β 267702 (2022).
@@
- KarabassovΒ etΒ al. [2022] T.Β Karabassov, I.Β V.Β Bobkova, A.Β A.Β Golubov,Β andΒ A.Β S.Β Vasenko,Β Hybrid helical state and superconducting diode effect in superconductor/ferromagnet/topological insulator heterostructures,Β Phys. Rev. BΒ 106,Β 224509 (2022).